A piece of sculpture is purchased for $3000 in 1960. In 2008,the sculpture is valued at $186000. What is the average annual rate of appreciation of the sculpture?
Answer by checkley71(8403)

186,000=3000(1+r)^48
186,000=3000(1+.089787)^48
186,000=3000(62)
186,000=186,000
Answer is 8.9787%.